### v1

# FaceEditTalker: Interactive Talking Head Generation with Facial Attribute Editing

Abstract
Recent advances in audio-driven talking head generation have achieved impressive results in lip synchronization and emotional expression. However, they largely overlook the crucial task of facial attribute editing. This capability is crucial for achieving deep personalization and expanding the range of practical applications, including user-tailored digital avatars, engaging online education content, and brand-specific digital customer service. In these key domains, the flexible adjustment of visual attributes—such as hairstyle, accessories, and subtle facial features—is essential for aligning with user preferences, reflecting diverse brand identities, and adapting to varying contextual demands. In this paper, we present FaceEditTalker, a unified framework that enables controllable facial attribute manipulation while generating high-quality, audio-synchronized talking head videos. Our method consists of two key components: an image feature space editing module, which extracts semantic and detail features and allows flexible control over attributes like expression, hairstyle, and accessories; and an audio-driven video generation module, which fuses these edited features with audio-guided facial landmarks to drive a diffusion-based generator. This design ensures temporal coherence, visual fidelity, and identity preservation across frames. Extensive experiments on public datasets demonstrate that our method outperforms state-of-the-art approaches in lip-sync accuracy, video quality, and attribute controllability. Project page: https://peterfanfan.github.io/FaceEditTalker/
1 Introduction
Figure 1: By providing a single reference image, audio input, and optional facial attribute input, our method generates high-quality, facially editable speaker videos by predicting facial landmark maps and performing linear edits on the feature semantic encoding of the image, combined with a diffusion model. This method demonstrates good generalization ability and achieves high lip-sync accuracy. In this figure, the image input used is a portrait from outside the dataset.
In recent years, audio-driven talking head generation  [38, 35, 53, 10, 16, 45, 30] has made remarkable progress and found widespread applications in domains such as virtual reality[28, 20], animation production[25, 51], online education[20], digital humans[16], and film post-production[28]. These technologies enable virtual characters to exhibit more natural and realistic speaking behaviors by synchronizing facial movements with audio input. However, most existing approaches primarily focus on lip synchronization [38, 35, 53, 10] and emotional expression [14, 50, 32, 13], while largely overlooking the important functionality of facial attribute editing.
Facial attribute editing is essential for audio-driven video generation due to its strong practical relevance. Beyond accurate audio-visual synchronization, users often require flexible control over visual appearance, including expressions, hairstyles, age, gender, and accessories like glasses. For example, virtual idols may need to adapt to different audience preferences, and digital customer service agents may need to reflect distinct brand identities. Dynamic and fine-grained attribute control can greatly enhance personalization and user engagement.
Although image-level facial attribute editing methods such as GAN-based or text-driven approaches [21, 23] have achieved initial success in static image generation tasks, extending these methods to video generation, particularly in talking-head videos, remains challenging. The main obstacles include maintaining temporal consistency of facial attributes, where edits must not only preserve realism within individual frames but also ensure smooth transitions across frames to avoid visual flickering or discontinuities, which is essential for producing natural video sequences [56, 49, 24]. Moreover, preserving audio-driven facial dynamics during attribute manipulation is crucial, as any inconsistency can disrupt accurate lip synchronization and natural facial motion.
To solve these problems, we propose FaceEditTalker, a novel method that seamlessly integrates audio-driven facial animation with controllable facial attribute editing. Our approach enables users to flexibly adjust facial attributes during video generation, while preserving high image fidelity, smooth facial dynamics, and accurate audio-visual alignment.
Our main contributions are summarized as follows:
(1) We propose a novel framework named FaceEditTalker, which seamlessly unifies facial attribute editing and audio-driven talking head generation. This framework enables fine-grained manipulation of attributes such as hair feature, facial structure, and accessories, while maintaining high-quality lip synchronization and natural motion dynamics.
(2) We propose an innovative two-stage heterogeneous latent diffusion model to address challenges in editability and consistency. Through editable latent feature extraction and feature re-injection, it enables highly flexible zero-shot editing while preserving identity integrity and natural dynamics.
(3) We conduct extensive experiments on multiple public datasets such as HDTF[60] and VoxCeleb2[33], demonstrating that our method significantly outperforms state-of-the-art baselines in terms of video quality, lip synchronization accuracy, keypoint alignment error, and identity consistency.
2 Related Work
Audio-driven Talking Head Generation. Recent methods for audio-driven talking head generation have made remarkable progress, emphasizing realism, identity preservation, and expression diversity. Early approaches [38, 35, 53, 10] primarily adopt encoder-decoder architectures to map audio signals to lip movements. While effective to some extent, these methods often suffer from blurred textures and weak identity preservation due to limited fusion strategies. To enhance realism, NeRF-based methods [16, 45, 30, 31, 37] and 3D Gaussian Splatting methods [2, 57, 29, 15] both model 3D geometry for more lifelike appearances; however, they typically require long video sequences and come with high computational costs. Another line of work leverages facial landmarks or 3D priors [61, 7, 59, 58] to disentangle speech content from identity features, improving controllability but often sacrificing fine-grained details in regions like lips and teeth. Recently, diffusion-based models [46, 48, 54, 44, 41, 6, 55, 17, 18, 2] have emerged as powerful solutions for high-quality and diverse talking head generation, with many of these approaches also leveraging facial landmarks as driving signals to enhance controllability and expression alignment. In our method, facial landmarks are adopted as controllable priors to guide a diffusion-based generator, achieving precise lip synchronization and identity preservation while supporting flexible facial attribute editing without compromising speech-driven facial dynamics.
Facial Attribute Editing. Facial attribute editing aims to manipulate specific attributes (e.g., age, hairstyle, glasses) while preserving identity. StyleGAN-based methods [21, 23, 43, 1] leverage latent space disentanglement to enable controllable editing, while CLIP-guided approaches [42, 36] introduce semantic alignment between language and image, allowing intuitive text-driven attribute modifications. Diffusion models offer enhanced control and fewer artifacts for editing tasks [5, 11, 40], but extending them to videos introduces temporal consistency challenges due to frame-wise randomness. To address this, methods like Latent-Transformer [56], STIT [49], and Diffusion-Video-Autoencoders [24] decompose identity and motion or apply global-local refinements to preserve coherence across frames. In our method, we perform linear transformations in the diffusion latent space while keeping the first frame fixed, ensuring both high-quality attribute editing and consistent facial dynamics across video frames.
3 Method
Figure 2: Overview of the inference process of our proposed framework FaceEditTalker. The framework consists of two main modules: (a) Image Feature Space Editing Module, which extracts editable semantic and stochastic codes from the reference image using a dual-layer latent encoding structure. Fine-grained attribute manipulation is enabled through optional spatial editing on the semantic codes. (b) Audio-Driven Video Generation Module, which leverages the audio input to infer driving landmarks. During the diffusion process, the stochastic codes guide dynamic generation, while the semantic codes serve as conditional inputs to ensure attribute consistency and visual fidelity throughout the video. The training procedure is detailed in Section 3.5 and Appendix A.
3.1 Overview
Our proposed framework, FaceEditTalker, introduces an innovative two-stage heterogeneous latent diffusion model to address challenges in editability and consistency. It consists of two tightly coupled modules: the Image Feature Space Editing Module and the Audio-Driven Video Generation Module. The Image Feature Space Editing Module extracts editable semantic and stochastic codes from the reference image using a dual-layer latent encoding structure, enabling fine-grained control over facial attributes, which can be further edited using text-guided linear classifiers. These features are then passed to the Audio-Driven Video Generation Module, where synchronized audio-driven landmarks guide a diffusion-based generative process to produce high-quality, temporally coherent talking head videos with consistent identity and natural lip-sync.
3.2 Task Formulation
[…]
3.3 Image Feature Space Editing Module
To achieve effective facial attribute editing, the Image Feature Space Editing Module leverages the design of DiffAE [40] with a dual-layer latent encoding structure. Inspired by the style vector mechanism in StyleGAN [21], our model decouples the latent space into two subspaces: semantic code $z_{\text{sem}}$ and stochastic code $Z_{T}$, capturing high-level semantic features and fine-grained details, respectively. This decomposition improves facial reconstruction accuracy and enhances controllability in attribute manipulation, supporting both zero-shot editing and fine-grained semantic control.
The semantic encoder $E_{\text{sem}}$ extracts global facial semantics from the input image, encoding them into low-dimensional vectors akin to StyleGAN’s style vectors, enabling linear transformations for attribute editing. Given a semantic code $z_{\text{sem}}$ and an attribute direction vector $w_{\text{attr}}$, attribute manipulation is expressed as:
$z^{\prime}_{\text{sem}}=z_{\text{sem}}+\alpha\cdot w_{\text{attr}},$ (1)
where $\alpha$ controls the intensity of the change. To enable text-driven editing and align more clearly with the illustrated process, text input is processed in conjunction with an attribute library by a linear classifier to compute the attribute direction vector $w_{\text{attr}}$. This $w_{\text{attr}}$ is then added to the original semantic code $z_{\text{sem}}$ according to Eq (1) to obtain the edited semantic code $z^{\prime}_{\text{sem}}$. A reconstruction loss $L_{\text{rec}}$ ensures that non-target regions remain consistent:
$L_{\text{rec}}=|z^{\prime}-z_{\text{static}}|.$ (2)
The stochastic encoder, designed as a Unet-based diffusion model [47]. First use the $E_{img}$ encode the input image into an initial latent representation $Z_{0}$. This representation undergoes a noise diffusion process to capture fine details, resulting in the final latent representation $Z_{T}$. We refer to this complete encoding and diffusion procedure as the stochastic encoder. The forward process, conditioned on $z_{\text{sem}}$, is defined as:
$z_{t+1}=\sqrt{\alpha_{t+1}}f_{\theta}(z_{t},t,z_{\text{sem}})+\sqrt{1-\alpha_{t+1}}\epsilon_{\theta}(z_{t},t,z_{\text{sem}}).$ (3)
This reverse diffusion effectively complements the semantic code’s lack of detail. During video generation, the semantic code $z_{\text{sem}}$ manages global attribute control, while the stochastic code $Z_{T}$ preserves local detail consistency, ensuring smooth and realistic facial transformations across frames.
3.4 Audio-Driven Video Generation Module
This module generates video using audio input, semantic code, and stochastic code. The audio input is first processed to extract audio features $A^{1:T}$. These features are then transformed into an audio-driven facial landmark sequence $L^{1:T}$ through a landmark prediction network, represented as:
$L^{1:T}:(l^{1},l^{2},\dots,l^{T})=P_{ldm}(A^{1:T}:(a^{1},a^{2},\dots,a^{T})),$ (4)
where $P_{ldm}$ encapsulates the transformation from audio features to landmarks, involving a pre-trained wav2vec model and a regression model.
Subsequently, the facial landmark sequence $L^{1:T}$ is processed by a landmark feature extractor to obtain the corresponding landmark features $F^{1:T}$:
$F^{1:T}:(f^{1},f^{2},\dots,f^{T})=E_{ldm}(L^{1:T}:(l^{1},l^{2},\dots,l^{T})),$ (5)
where $E_{ldm}$ represents the landmark feature extractor. It adopts multi-scale strategies and cross-attention mechanisms to capture multi-level dynamics and enhance feature correlations, improving precision and temporal consistency in video generation.
The extracted facial landmark features $F^{1:T}$ is fused with the stochastic code $Z_{T}$ to produce the feature $K$. During the sampling process of the diffusion model, we employ an innovative conditional feature injection strategy, where the high-level semantic information $z_{\text{sem}}$ is injected as a global facial attribute control signal into the conditional diffusion model, while the target feature $K$ serves as dynamic motion information. This mechanism ensures stable expression of semantic attributes alongside synchronized audio-driven facial movements:
$z_{t-1}=\sqrt{\alpha_{t}}z_{t}+\sqrt{1-\alpha_{t}}\epsilon_{\theta}(z_{t},z_{\text{sem}},K,t),$ (6)
where $\alpha_{t}$ represents the noise scheduling parameters, and $\epsilon_{\theta}$ is the conditional denoising network that guides the denoising process based on the latent variable $z_{\text{sem}}$ and the target features $K$.
Finally, the denoised latent variable $z_{0}$ is decoded into a sequence of video frames, which not only exhibit dynamic facial expressions synchronized with the audio, but also, depending on the switch selection, achieve facial attribute editing of the speaker based on the different semantic encoding $z_{\text{sem}}$, resulting in a high-quality editable talking head video.
3.5 Training and Inference Pipeline
Training process: This model is divided into three stages. First stage jointly trains the semantic encoder and random encoder to extract semantic and random detail features of images by minimizing the mean squared error between predicted and actual noise, with the loss function:
$\mathcal{L}_{\text{simple}}=\sum_{t=1}^{T}\mathbb{E}_{x_{o},\epsilon_{t}}[||\epsilon_{\theta}(x_{t},t,z_{\text{sem}})-\epsilon_{t}||_{2}^{2}].$ (7)
Second stage trains an image semantic linear classifier, optimizing the model using cross-entropy loss to accurately classify image attributes, with the loss function:
$\mathcal{L}=-\frac{1}{N}\sum_{l=1}^{N}[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})].$ (8)
Third stage trains the latent space diffusion model generation network, fusing semantic and key point features, and generating images by minimizing the mean squared error between predicted and actual noise, with the loss function:
$\mathcal{L}=\sum_{t=1}^{T}\mathbb{E}_{z_{o},\epsilon_{t}}[||\epsilon_{\theta}(z_{t},t,z_{\text{sem}},z_{l})-\epsilon_{t}||_{2}^{2}].$ (9)
Inference process: First, face key points and pose information are extracted from the audio sequence and reference image, then combines semantic encoding and attribute editing information and inputs them into the trained latent space diffusion model to finally generate the video sequence of a talking head.
4 Experiments
4.1 Experimental Settings
Datasets. We trained the encoder of the dual-layer latent architecture using the FFHQ dataset [22], which offers high-resolution facial images with diverse attributes including age, race, expression, facial structure, hair features, and accessories, ideal for learning complex feature representations. For the linear classifier, we used the CelebA-HQ dataset [26] with binary labels for 40 facial attributes to enhance attribute feature separation and model generalization. In the audio-driven facial animation generation stage, we utilized the HDTF dataset [60], containing lip-sync videos from over 300 speakers, along with VoxCeleb2 [33] and VFHQ [52] datasets to improve the model’s ability to learn complex mappings between speech and facial movements under various environmental conditions. Additionally, we applied LatentSync [27] to refine dataset quality by resampling videos, removing those with low synchronization confidence, correcting audiovisual offsets, and filtering out clips with poor HyperIQA scores, thereby enhancing lip-sync accuracy and visual quality.
Comparison Methods. To the best of our knowledge, there is no existing method capable of generating high-resolution, audio-driven speaker videos with editable facial attributes. For a comprehensive evaluation of our proposed method, we first generate results using semantic features extracted by the high-level semantic encoding module, ensuring identity consistency with reference images. Our method is compared against several state-of-the-art (SOTA) lip synchronization approaches categorized into three groups. Wav2Lip [38] optimizes direct mappings between audio and lip motion for highly synchronized lip movements while preserving facial textures. SadTalker [58] employs explicit facial landmarks and adversarial networks to produce smooth animations. DiffTalk [46], EchoMimic [9], and Hallo [54] leverage diffusion models to model conditional distributions between audio and facial movements, achieving higher-quality talking videos and strong generalization capabilities for out-of-distribution subjects. This comparison aims to evaluate our method’s performance relative to current leading techniques in audio-driven talking head generation.
Evaluation Metrics. For evaluating our method, we employ several metrics. Image generation quality is assessed using FID [12], SSIM [3], PSNR [19], and CPBD [34]. Lip motion accuracy is evaluated with M-LMD and F-LMD [8], while Syncconf [39] measures lip movement-audio synchronization. Additionally, we edit semantic features from the dual-layer semantic encoding module using a linear classifier to produce edited video results. These are compared against state-of-the-art video editing methods such as Latent-Transformer [56], STIT [49], and Diffusion-Video-Autoencoders [24], using TL-ID and TG-ID [49] as evaluation metrics.
4.2 Comparison with Other Methods
Table 1: Quantitative evaluation of our approach compared with SOTAs. Our method achieves the best performance on FID, SSIM (second best in EchoMimic), lip sync metrics, and keypoint error metrics, demonstrating superior overall quality and synchronization accuracy.
Video Quality Lip Sync Keypoint Error
Method FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ CPBD $\uparrow$ Min Dist $\downarrow$ AVConf $\uparrow$ AVOffset($\rightarrow 0$) M-LMD $\downarrow$ F-LMD $\downarrow$
Real Video (HDTF) 0.000 1.000 35.668 0.263 7.238 8.993 0.000 0 0
Wav2Lip[38] 20.641 0.532 16.929 0.199 6.611 8.119 -2.000 4.368 4.256
SadTalker[58] 25.566 0.698 22.211 0.204 8.527 3.163 1.000 3.368 3.192
DiffTalk[46] 18.570 0.558 26.587 0.225 10.091 3.046 -4.000 5.473 1.146
EchoMimic[9] 17.486 0.893 25.968 0.210 9.163 6.146 -1.000 3.983 3.790
Hallo[54] 16.880 0.821 25.331 0.203 9.612 6.128 0.000 3.412 3.532
Our Method 16.580 0.843 25.574 0.205 9.527 6.354 0.000 3.354 3.465
Real Video (VoxCeleb2) 0.000 1.000 26.453 0.272 7.701 6.365 0.000 0 0
Wav2Lip[38] 20.565 0.468 16.042 0.201 7.665 8.236 -2.000 4.368 4.256
SadTalker[58] 23.421 0.634 21.254 0.211 13.542 3.355 1.000 3.368 3.192
EchoMimic[9] 17.586 0.910 24.948 0.209 9.654 6.542 -1.000 3.983 3.790
Hallo[54] 15.785 0.751 25.738 0.188 8.142 6.105 0.000 3.408 3.498
Our Method 15.418 0.772 25.985 0.189 8.068 6.252 0.000 3.354 3.465
Quantitative Evaluation. We quantitatively compared FaceEditTalker with existing state-of-the-art audio-to-face generation methods on the HDTF and VoxCeleb2 datasets. As shown in the table, FaceEditTalker demonstrates excellent performance in metrics like image feature similarity, structural similarity, and image quality, leveraging a latent space diffusion model. Our method achieves significant improvements compared to previous diffusion-based approaches, attributed to a more advanced model framework. Furthermore, our model shows strong SyncNet scores, benefiting from optimizations such as dataset correction, multi-scale strategies, and cross-attention mechanisms.
Table 2: Quantitative results of our approach compared with SOTAs. Our method achieves the best performance on identity consistency in face attribute editing.
Method TL-ID $\uparrow$ TG-ID $\uparrow$
Latent-Transformer[56] 0.975 0.913
STIT[49] 0.990 0.969
Diffusion-Video-Autoencoders[24] 0.986 0.991
Our Method 0.992 0.989
Furthermore, regarding identity consistency in face attribute editable talking head generation, we conducted quantitative evaluation. Our method edits semantic features to generate videos with 20 attributes, compared against video editing algorithms. Evaluation of identity consistency between frames showed that while our generative model achieved similar overall identity consistency as video editing methods, it significantly excelled in TL-ID.
Figure 3: Qualitative evaluation compared with other methods. Using two different reference images and the same audio clip, our method is tested without enabling the editing feature. Our approach demonstrates superior performance in both facial expression naturalness and video quality.
Qualitative Evaluation. Figure 3 compares the generation quality of FaceEditTalker with existing advanced methods. While previous approaches often prioritize specific aspects such as lip synchronization or introduce artifacts and distortions when aiming for expressiveness, our proposed algorithm generates videos with better overall image quality and more accurately captures fine facial expression details, closely matching the original video. FaceEditTalker particularly excels at handling nuanced facial actions like eye closure and mouth opening, contributing to a higher level of realism in the generated results.
Figure 4: Video generation results with the editing feature enabled. Using three different reference images and the same audio clip, we demonstrate the editing and speaker generation effects under four different attribute editing categories with various sub-attributes. More attribute editing results can be found in Appendix C and Supplementary materials.
Table 3: User Study Results for Editable Facial Attribute Talking Head Generation
Method Lip Sync $\uparrow$ Realism $\uparrow$ Video Quality $\uparrow$ Attribute Editing Effect $\uparrow$
Original Video 4.80 4.90 4.80 Not Supported
Wav2Lip[38] 3.60 3.10 3.70 Not Supported
SadTalker[58] 2.70 2.20 2.80 Not Supported
DiffTalk[46] 3.40 3.00 2.90 Not Supported
EchoMimic[9] 3.40 3.70 3.90 Not Supported
Hallo[54] 3.80 3.30 3.60 Not Supported
Latent-Transformer[56] 3.30 3.80 3.60 3.40
STIT[49] 3.30 3.30 3.00 3.30
Diffusion Video Autoencoders[24] 3.10 3.50 3.10 3.30
Our Method 3.80 3.60 3.70 4.40
User Study. We conducted a user study (10 people, score 1-5) on lip sync accuracy, realism, video quality and attribute editing effects using a five-point scale. Our method achieved high scores across all metrics: 3.8 for lip-sync (benefiting from multi-scale keypoint features and SyncNet preprocessing), strong performance in realism and video quality (attributed to the latent space diffusion model’s capabilities), and superior performance in attribute editing. As shown in Table 3, methods marked with ’N/A’ in the Attribute Editing Effect column indicate that they do not support this functionality.
4.3 Analysis and Ablation Study
Table 4: Quantitative Metrics for Ablation Study of Facial Landmark Feature Extractor
Method Min Dist $\downarrow$ AVConf $\uparrow$ AVOffset($\rightarrow 0$) M-LMD $\downarrow$ F-LMD $\downarrow$
Original Video 7.359 7.586 0.000 0 0
Multi-layer Convolution 12.534 6.562 7.000 10.468 7.892
Multi-scale Strategy 9.300 5.344 3.000 4.532 5.246
Multi-scale & Cross-Attention 8.145 6.288 0.000 3.354 3.465
[…]
Effectiveness of the Facial Landmark Feature Extractor. We conducted ablation experiments on two datasets to validate the multi-scale strategy and cross-attention mechanism in the facial landmark feature extractor. The multi-scale strategy improved fine-grained movements like lip and eye motions, while the cross-attention mechanism enhanced all metrics, particularly AVOffset and M-LMD, achieving better consistency with the original video. Furthermore, testing the Edited Semantic Encoding showed minimal impact on evaluation metrics.
Figure 5: Principal Component Analysis (PCA) visualization of the four attributes.
Figure 6: Talking head generation results after interpolating in the high-level semantic feature space using two reference images.
Linear Distribution of Attributes in Latent Space. To verify the linear distribution of target attributes in latent space, we first visualized the latent space using Principal Component Analysis (PCA). The results show that in the PCA space, samples of different attributes exhibit clear separation along the principal component directions, proving that semantic latent variables exhibit a linear distribution in the latent space. This validates the rationality of linear operations in the semantic space and supports the classifier’s ability to distinguish between different attribute categories. Additionally, to demonstrate the interpolability of high-level semantic features across different attributes, we generated videos by interpolating features from two different speakers’ images. The interpolation results appeared very natural, showing good separation of attributes.
5 Conclusion and Limitations
Conclusion. In summary, we have introduced a novel framework for editable talking face generation that significantly enhances both the realism and controllability of facial animations. By leveraging a combination of disentangled latent representations and fine-grained audio-visual alignment, our method enables intuitive editing capabilities such as face editing and lip-sync correction. Extensive experiments on multiple benchmarks demonstrate that our approach not only achieves superior performance compared to existing state-of-the-art methods but also allows for diverse and personalized talking face generation. We believe this framework opens up promising directions for future research in personalized avatars, virtual assistants, and digital human synthesis.
Limitations. Despite the success of our framework, we also recognize some limitations during the exploration. First, while demonstrating good generalization ability, performance might degrade in environments with highly diverse identities, complex head movements, or challenging conditions that are not well-represented in the training data. Second, the diffusion-based generation process, although yielding high-quality results, currently incurs significant computational costs, limiting its application in real-time scenarios. Third, currently our method can only support the editing of 40 preset attributes. In the future, we hope to leverage the capabilities of models like CLIP to achieve flexible face attribute editing through arbitrary text descriptions.
Ethical Considerations. Our talking head generation method is strictly limited to academic research and will not be applied to unethical domains like fraud or misinformation. The generated model will be shared with the Deepfake detection community to support identification research, ensuring responsible development of this technology.
6 Acknowledgments
The work was jointly supported by the National Science and Technology Major Project under grant No. 2022ZD0117103, the National Natural Science Foundations of China under grant No. 62272364, 62472342, the Provincial Key Research and Development Program of Shaanxi under grant No. 2024GH-ZDXM-47, the Research Project on Higher Education Teaching Reform of Shaanxi Province under Grant No. 23JG003.
Appendix A Pseudo-code for this method
A.1 Training and Inference Process Description
The training phase consists of three main modules:
1. Joint Training of Semantic Encoder and Random Encoder: Face images are encoded to extract high-level semantic and random features, optimized using a diffusion model.
2. Training the Image Semantic Linear Classifier: Attribute variation directions are learned in the high-level semantic space, enabling precise attribute classification.
3. Training the Latent Space Diffusion Model: Multi-scale facial motion features, semantic encodings of reference images, and audio-driven information are integrated to generate high-quality, identity-consistent talking face videos.
In the inference phase, the model uses the input audio sequence, reference image, and attribute information to generate facial motion features, which are processed by the diffusion model to produce high-quality, editable talking face videos.
A.2 Training Stage
Algorithm 1 Joint Training of Semantic Encoder and Random Encoder
1:Input: Face image set (with attribute labels), attribute list, learning rate, diffusion time step
2:Output: Weight vectors corresponding to each attribute list
3:for each epoch do
[…]
5:         $z_{\text{sem}}=\text{SemanticEncode}(x_{i})$triangleright Semantic Encoder Forward Pass
6:         $t=\text{RandomTimeStep}()$
7:         $z_{0}=\text{Encode}(x_{i})$
8:         $x_{t}=\text{AddNoise}(z_{0})$
9:         $\hat{x}=\hat{q}(x_{t},z_{\text{sem}})$triangleright Model Forward Pass
10:         $e=\text{ComputeTarget}(z_{t})$triangleright Compute Target
11:         $L=\text{MSE}(e,\hat{x})$triangleright Loss Calculation
12:         Backpropagate and update parameters
13:     end for
14:     Save model at the end of each epoch
15:end for
Algorithm 2 Training the Image Semantic Linear Classifier
1:Input: Face image set, attribute labels, learning rate
2:Output: Weight vectors corresponding to each attribute list
3:for each epoch do
4:     for each attribute $a_{i}\in\text{Attr}(a_{1},a_{2},\dots,a_{n})$ do
5:         $z_{\text{sem}}=\text{SemanticEncode}(x_{i})$
6:         $y_{\text{labels}}=\text{GetAttributeLabels}(a_{i},\text{FFHQ}(image))$
[…]
8:         $b=\text{InitializeBias}()$
[…]
10:              for each $z_{\text{sem}}$ do
11:                  $y_{\hat{t}}=\text{Sigmoid}(W^{T}z_{\text{sem}}+b)$triangleright Forward Pass
12:                  $L=\text{CrossEntropyLoss}(y_{\text{labels}},y_{\hat{t}})$triangleright Cross-Entropy Loss
13:                  Backpropagate and update parameters
14:              end for
15:         end for
16:     end for
17:     Save weight vectors and attribute-label pairs
18:end for
Algorithm 3 Training the Latent Space Diffusion Model Generation Network
1:Input: Face image set, reference image set, learning rate
2:Output: Diffusion model parameters
3:for each epoch do
4:     for each batch $x_{r},x_{\text{ref}},l_{r},l_{\text{ref}}$ do
5:         $z_{0}=\text{Encode}(x_{r})$
6:         $z_{t}=\text{AddNoise}(z_{0})$
7:         $t=\text{RandomTimeStep}()$
8:         $z_{\text{ref}}=\text{SemanticEncode}(x_{\text{ref}})$
9:         $z_{1}=\text{PoseGuided}(l_{r},l_{\text{ref}})$
10:         $y_{\hat{t}}=\hat{q}(z_{1},z_{t},z_{\text{ref}})$
11:         $e=\text{ComputeTarget}(z_{t})$
12:         $L=\text{MSE}(e,y_{\hat{t}})$triangleright MSE Loss
13:         Backpropagate and update parameters
14:     end for
15:     Save model at the end of each epoch
16:end for
A.3 Inference Stage
Algorithm 4 Based on Linear Space Facial High-Level Semantic Feature Editing Module (Inference Process)
1:Input: Reference image $x_{ref}$, audio sequence $\mathcal{S}_{audio}$, attribute labels $(y,w)$, attribute editing magnitude, diffusion time steps steps
2:Output: Speaker video frame sequence $S_{video}$
3:Data Preprocessing
4:Use Wav2Vec to extract audio feature sequence from audio sequence $\mathcal{S}_{audio}$
5:$F^{1:T}=F(f_{1},f_{2},...,f_{t})=Wav2Vec(S_{audio})$
6:$x_{mesh}=MediaPipeMesh3D(x_{ref})$
7:$M^{1:T}=Audio2Mesh(F^{1:T},x_{mesh})$
8:$P^{1:T}=Audio2Pose(F^{1:T})$
9:$L^{1:T}=PerspectiveProjection(M^{1:T},P^{1:T})$
10:$l_{ref}=MediaPipeFaceLandmarker(x_{ref})$
11:$z_{l}^{1:T}=PoseGuider(L^{1:T},l_{ref})$
12:$z_{sem}=SemanticEncode(x_{ref})$
13:if a is not None then
14:     $z_{sem}=z_{sem}+\alpha w$
15:end if
16:$S_{video}=DiffusionModel(z_{sem},steps,z_{l}^{1:T})$
Appendix B Experimental parameter settings
Our methods was trained on two A100 GPUs. The first and second stages were trained for 100 hours, and the third stage was trained for 160 hours. The main parameters are shown in Table 5.
Table 5: Experimental Parameter Settings
Parameters Value/Range
Random Seed 0
Image Size 512*512
Batch Size 16
Learning Rate 0.0001
Training Epochs 20000
Embedding Layer Channels 512
Diffusion Timesteps 1000
Appendix C More Qualitative Evaluation
More qualitative evaluation can be seen in Figures 7-14, which include 40 attribute editing effects of two celebrities, and the dynamic video can be found in the attachment.
Figure 7: Accessories & Makeup of Jay Chou
Figure 8: Accessories & Makeup of Michelle Yeoh
Figure 9: Facial State of Jay Chou
Figure 10: Facial State of Michelle Yeoh
Figure 11: Facial Structure of Jay Chou
Figure 12: Facial Structure of Michelle Yeoh
Figure 13: Hair Features of Jay Chou
Figure 14: Hair Features of Michelle Yeoh